OpenStudy (anonymous):

If x varies inversely as y and directly as t, and x = 12 when t = 10 and y = 25, find y when x is 6 and t = 3. 9 3/5 15 166 2/3

OpenStudy (anonymous):

The first step is to create your model.

OpenStudy (anonymous):

Looks like x = k*t*(1/y)

OpenStudy (anonymous):

You are given values for x, t, and y so that you can solve for k.

OpenStudy (anonymous):

Are you with me?

OpenStudy (anonymous):

not sure....

OpenStudy (anonymous):

What is giving you difficulty?

OpenStudy (anonymous):

the solve for k....how should i do it???

OpenStudy (anonymous):

You are given a word problem that you need to use to create a model. Your problem says x increases directly with t and inversely with y. Start with a general equation that grows directly with t and inversely (1/y) with y.

OpenStudy (anonymous):

The general equation described above is x = k*t*(1/y) This simple expression is useful because it generally describes all situations where x and y are inversely related and x and t are directly related. Your question goes into more detail giving you values of x, y, and t so that you can solve for the constant factor k.

OpenStudy (anonymous):

Step one is to setup the model \[x = kt(\frac{ 1 }{ y })\] Step 2 is to plug in x = 12 when t = 10 and y = 25 and solve for k. Are you with me?

OpenStudy (anonymous):

ok like this....12=k*10(1/25)

OpenStudy (anonymous):

Yes! :-)

OpenStudy (anonymous):

but what is the first step to solve for k???

OpenStudy (anonymous):

use arithmetic to get k isolated so that you have # = k

OpenStudy (anonymous):

I guess the first step would be to multiply both sides by 25.

OpenStudy (anonymous):

like this...25*12=k*10*1/25*25 25*12=k*10

OpenStudy (anonymous):

right

OpenStudy (anonymous):

what should i do next???

OpenStudy (anonymous):

divide both sides by 10

OpenStudy (anonymous):

which =30????

OpenStudy (anonymous):

Very good. Now step 3 is to take the second set of x's t's and (along with k) use the model equation to solve for y.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

gat it is it 15???

OpenStudy (anonymous):

:-D Nice job!

OpenStudy (anonymous):

thankx for ur help and time...:)

OpenStudy (anonymous):

You are welcome. Study hard!